You purchased a zero-coupon bond one year ago for $277.83. The market interest rate is now 9 percent. Assume semiannual coumpounding periods.


If the bond had 15 years to maturity when you originally purchased it, what was your total return for the past year?
Answer #1

Price of the bond 14 years to maturity:-

=PV(rate,nper,pmt,fv)

=PV(9%/2,14*2,0,1000)

=291.57

Total return =291.57-277.83 =13.74$

In % return =291.57/277.83-1 =4.95%